# Riemannian Variational Flow Matching for Material and Protein Design

### v2

@@ Results.
We observe that (1) Riemannian models better capture manifold geometry by generating points with minimal distance to the manifold compared to Euclidean ones (see Distance columns in table 1), and (2) variational models produce sharper and less blurred distributions than vanilla models, with RG-VFM-$\mathbb{R}^{3}$ and RG-VFM-$\mathcal{M}$ showing the best visual performance in fig. 3. This is reflected in table 1 in Coverage metric results, where variational models – particularly Riemannian ones – achieve the highest values. For C2ST, no consistent pattern emerges between spherical and hyperbolic cases, except that in both cases standard VFM demonstrates the strongest performance. In essence, emphasizing endpoint accuracy enables variational models to capture fine details of the target distribution’s shape, and additional geometric awareness of RG-VFM further enhances the result. Additionally, preliminary findings suggest that using $L^{1}$ loss instead of $L^{2}$ may enhance performance, particularly in hyperbolic spaces, and we start exploring this option in section E.4.
@@ 5.3 Protein Backbone Generation with V-ReQFlow
QFlow (Yue et al., 2025) is a flow-based model for protein backbone generation. Unlike previous methods (Yim et al., 2023b; Bose et al., 2023) that represent $\mathrm{SO}(3)$ elements with rotation matrices, QFlow uses quaternions, which provide improved training stability. Building on this foundation, ReQFlow (Yue et al., 2025) further enhances QFlow by incorporating rectified flow with re-paired samples and noise, inspired by Liu et al. (2023), improving the designability of generated protein backbone structures. Similar to MOF structure generation, protein backbone structures are represented as sequences of $SE(3)$ elements $\{q_{i},t_{i}\}^{N}$, where $q_{i}\in\mathrm{SO}(3)$ defines the frame on the $\alpha$-carbon of each amino acid, and $t_{i}\in\mathbb{R}^{3}$ represents the zero-mean coordinate of the $\alpha$-carbon. For further details, see Yim et al. (2023b). The goal is to learn a conditional flow $p_{\theta}(Q,T|N)$ where $Q=\{q_{i}\}^{N}$ and $T=\{t_{i}\}^{N}$, with $N$ denoting the number of residues in the desired backbone structure. Like MOFFlow, both QFlow and ReQFlow employ a re-parametrized training objective that predicts end points from which vector fields are reconstructed. We apply our method by variationalizing the rotational component of their loss function, similarly to section F.1, while maintaining all other implementation details identical to isolate the benefits of our variational objective.
@@ D.4 A Comparison based on Jacobi Fields
For $s=0$, we are interested in the geodesic $\gamma_{0}:\tau\to\text{exp}_{x_{0}}(\tau v^{0})$, with $v^{0}$ the target velocity, and for $s=1$, the geodesic $\gamma_{1}:\tau\to\text{exp}_{x_{0}}(\tau v^{1})$ is defined with $v^{1}$ the velocity learned by the model. $w=v^{0}-v^{1}$ corresponds to the discrepancy between the learned and the conditional initial velocity, and their norm is exactly what is minimized in the vanilla Riemannian Flow Matching (Chen & Lipman, 2024).
@@ F.1 Experimental Setup
which corresponds to the squared geodesic distance between $q_{1}$ and $\hat{q}_{1}$ in $SO(3)$. The definition of the vector fields is unchanged from eqs. 28 and 29, as well as the sampling algorithm.

### v3

@@ Results.
We observe that (1) Riemannian models better capture manifold geometry by generating points with minimal distance to the manifold compared to Euclidean ones (see Distance columns in table 1), and (2) variational models produce sharper and less blurred distributions than vanilla models, with RG-VFM-$\mathbb{R}^{3}$ and RG-VFM-$\mathcal{M}$ showing the best visual performance in fig. 3. This is reflected in table 1 in Coverage metric results, where variational models – particularly Riemannian ones – achieve the highest values. For C2ST, no consistent pattern emerges between spherical and hyperbolic cases, except that in both cases standard VFM demonstrates the strongest performance. In essence, emphasizing endpoint accuracy enables variational models to capture fine details of the target distribution’s shape, and additional geometric awareness of RG-VFM further enhances the result. We tested setting $\sigma_{t}(x)=1$ versus $\sigma_{t}(x)=1-t$ in eq. 8 during training and found negligible differences, so we report only the results obtained with $\sigma_{t}(x)=1$. Additionally, preliminary findings suggest that using $L^{1}$ loss (which corresponds to using a Riemannian Laplace instead of a Riemannian Gaussian distribution in eq. 8) instead of $L^{2}$ may enhance performance, particularly in hyperbolic spaces, and we start exploring this option in section E.4.
@@ 5.3 Protein Backbone Generation with V-ReQFlow
QFlow (Yue et al., 2025) is a flow-based model for protein backbone generation. Unlike previous methods (Yim et al., 2023b; Bose et al., 2023) that represent SO(3){SO}(3) elements with rotation matrices, QFlow uses quaternions, which provide improved training stability. Building on this foundation, ReQFlow (Yue et al., 2025) further enhances QFlow by incorporating rectified flow with re-paired samples and noise, inspired by Liu et al. (2023), improving the designability of generated protein backbone structures. Similar to MOF structure generation, protein backbone structures are represented as sequences of $SE(3)$ elements $\{q_{i},t_{i}\}^{N}$, where $q_{i}\in{SO}(3)$ defines the frame on the $\alpha$-carbon of each amino acid, and $t_{i}\in\mathbb{R}^{3}$ represents the zero-mean coordinate of the $\alpha$-carbon. For further details, see Yim et al. (2023b). The goal is to learn a conditional flow $p_{\theta}(Q,T|N)$ where $Q=\{q_{i}\}^{N}$ and $T=\{t_{i}\}^{N}$, with $N$ denoting the number of residues in the desired backbone structure. Like MOFFlow, both QFlow and ReQFlow employ a re-parametrized training objective that predicts endpoints from which vector fields are reconstructed. We apply our method by variationalizing the rotational component of their loss function, similarly to section F.1, while maintaining all other implementation details identical to isolate the benefits of our variational objective.
@@ D.2 How does RG-VFM fit in the Existing Flow Matching Framework?
Algorithm 3 CFM
0:  base $p\in{\mathbb{R}^{d}}$, target $q\in\mathbb{R}^{d}$.
#\# Training Phase
 Initialize parameters $\theta$ of $v_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\norm{v_{t}(x_{t};\theta)-\dot{x}_{t}}_{g}^{2}}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
[…]
Algorithm 4 VFM
0:  base $p\in{\mathbb{R}^{d}}$, target $q\in\mathbb{R}^{d}$.
#\# Training Phase
 Initialize parameters $\theta$ of $\mu_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{0.5,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0}\pgfsys@color@cmyk@stroke{0}{0}{1}{.5}\pgfsys@color@cmyk@fill{0}{0}{1}{.5}\norm{\mu_{t}(x_{t};\theta)-x_{1}}_{g}^{2}}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
compute corresponding velocity:
$\color[rgb]{0.5,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0}\pgfsys@color@cmyk@stroke{0}{0}{1}{.5}\pgfsys@color@cmyk@fill{0}{0}{1}{.5}\dot{x}_{t}=\frac{\mu_{t}(x_{t};\theta)-x_{t}}{1-t}$}
[…]
Algorithm 5 RFM
0:  base $p\in{\mathcal{M}}$, target $q\in\mathcal{M}$.
#\# Training Phase
 Initialize parameters $\theta$ of $v_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\norm{v_{t}(x_{t};\theta)-\dot{x}_{t}}_{g}^{2}}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
[…]
Algorithm 6 RG-VFM (general)
0:  base $p$, target $q\in\mathcal{M}$.
#\# Training Phase
 Initialize parameters $\theta$ of $\mu_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{dist}_{g}^{2}(x_{1},\mu_{t}(x_{t};\theta))}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
compute corresponding velocity $\dot{x}_{t}$
[…]
Algorithm 7 RG-VFM intrinsic
0:  base $p\in{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\mathcal{M}}$, target $q\in\mathcal{M}$.
#\# Training Phase
 Initialize parameters $\theta$ of $\mu_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{dist}_{g}^{2}(x_{1},\mu_{t}(x_{t};\theta))\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
compute corresponding velocity:
$\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\dot{x}_{t}=\frac{\texttt{log}_{x_{t}}(\mu_{t}(x_{t};\theta))}{1-t}$}
[…]
Algorithm 8 RG-VFM extrinsic
0:  base $p\in{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbb{R}^{d}}$, target $q\in\mathcal{M}$.
#\# Training Phase
 Initialize parameters $\theta$ of $\mu_{t}$
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\text{dist}_{g}^{2}(x_{1},\mu_{t}(x_{t};\theta))\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
compute corresponding velocity:
$\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\dot{x}_{t}=\frac{\mu_{t}(x_{t};\theta)-x_{t}}{1-t}$}
[…]
@@ D.4 A Comparison based on Jacobi Fields
For $s=0$, we are interested in the geodesic $\gamma_{0}:\tau\to\text{exp}_{x_{0}}(\tau v^{0})$, with $v^{0}$ the target velocity, and for $s=1$, the geodesic $\gamma_{1}:\tau\to\text{exp}_{x_{0}}(\tau v^{1})$ is defined with $v^{1}$ the velocity learned by the model. $w=v^{0}-v^{1}$ corresponds to the discrepancy between the learned and the conditional initial velocity, and their norm is exactly what is minimized in the vanilla Riemannian Flow Matching (Chen and Lipman, 2024).
@@ F.1 Experimental Setup
which corresponds to the squared geodesic distance between $q_{1}$ and $\hat{q}_{1}$ in $SO(3)$. The definition of the vector fields is unchanged from eqs. 29 and 30, as well as the sampling algorithm.